# Optimal multiple-phase estimation with multi-mode NOON states against photon loss

### v1

@@ I Introduction
In this work, we propose an optimal multiple-phase estimation scheme using the weighted multi-mode NOON states in the presence of photon loss. Analytically deriving the estimation precision for the proposed scheme, we investigate the robustness of the scheme against photon loss and compare it with the scheme using a particular type of the multi-mode NOON state proposed in Ref. p.c.humphreys , which is known to be optimal in the lossless case. A quantum advantage is also studied in comparison with the SQL set by a scheme using the weighted multi-mode coherent states. As a result, we find that our scheme is not only more robust against photon loss than the scheme in Ref. p.c.humphreys , but also exhibits a quantum advantage over the SQL even in the presence of loss. These findings are elaborated in more details for the cases of three- and four-mode two-photon NOON states that have been experimentally demonstrated s.hong ; s.hong2 , followed by the generalization of the scheme with increasing the numbers of modes and photons. The proposed schemes are also investigated with a measurement scheme with general structure consisting of a multi-mode beam-splitter and multiple photon-number-resolving detectors (PNRDs), offering a practical means to achieve the quantum advantage in multiple-phase estimation under a lossy environment. We expect that our results and methodology can be used for various studies in robust quantum metrology, given that loss and decoherence are inevitable.
@@ II Multiple-phase estimation scheme under lossy environment
Figure 1: Multiple-phase estimation using a $d+1$-mode probe state for estimating $d$ unknown phases $\{\phi_{1},\cdots,\phi_{d}\}$ with a measurement. Without loss of generality, the mode 0 is considered as a reference. It is assumed that photon loss rate in the mode $j$ is $\gamma_{j}\in[0,1]$.
@@ II.1 Scheme
Let us consider a scheme for simultaneously estimating $d$ unknown phases $\bm{\phi}=\{\phi_{1},\cdots,\phi_{d}\}$ in the presence of photon loss as shown in Fig. 1. We assume that the mode 0 is used as a reference upon which the phases of all the other modes are defined. For the multiple-phase estimation, we employ a ($d+1$)-mode NOON state probe written by
@@
where $\hat{a}_{j}^{\dagger}$ is the photon creation operator in the mode $j$ and $p_{j}$ with $\sum_{j=0}^{d}p_{j}=1$ and $p_{j}\in[0,1]$ is the weight of the mode $j$. This probe state undergoes $d$ individual phase shifts described by a unitary operator $\hat{U}_{\bm{\phi}}=\bigotimes_{j=1}^{d}\exp(i\phi_{j}\hat{a}_{j}^{\dagger}\hat{a}_{j})$. For practical relevance, we take into account photon loss that could occur in the transmission channels or inefficient detectors. The photon loss can be modeled by a fictitious beam-splitter represented by the mode transformation $\hat{a}_{j}^{\dagger}\rightarrow\sqrt{1-\gamma_{j}}\hat{a}_{j}^{\dagger}+\sqrt{\gamma_{j}}\hat{e}_{j}^{\dagger}$ p.kok with photon loss rate $\gamma_{j}\in[0,1]$, where $\hat{e}_{j}^{\dagger}$ is a virtual vacuum input mode for describing environment. It can be shown, by the Kraus representation of photon loss, that photon losses occurring before and after the phase shifters are equivalent. This allows to consider only the photon loss taking place onto the phase-encoded probe state $|\psi_{N,\bm{\phi}}\rangle=\hat{U}_{\bm{\phi}}|\psi_{N}\rangle$ before measurement. The phase-encoded ($d+1$)-mode NOON state is thus transformed to the output state
$\hat{\rho}_{N,\bm{\phi}}=\Lambda_{\gamma_{0}}\otimes\cdots\otimes\Lambda_{\gamma_{d}}(|\psi_{N,\bm{\phi}}\rangle\langle\psi_{N,{\bm{\phi}}}|),$ (2)
where $\Lambda_{\gamma_{j}}$ is a photon loss channel in the mode $j$ m.grassl . If $\gamma_{j}$ is zero, then the quantum channel $\Lambda_{\gamma_{j}}$ becomes an identity channel. Finally, the estimation of $d$ unknown phases is made based on the measurement results obtained by a measurement performed onto the output state of Eq. (2). We also consider an estimation scheme with classical light to examine a quantum advantage of the proposed scheme through quantum parameter estimation theory that is introduced below. As a classical benchmark, we consider the scheme that employs a ($d+1$)-mode coherent state described by $\otimes_{j=0}^{d}|\alpha_{j}\rangle$ with $\alpha_{j}=\sqrt{q_{j}N}$ ($q_{0}+\cdots+q_{j}=1$, $q_{j}\in[0,1]$) under the same environment, leading to the output state $|\sqrt{1-\gamma_{0}}\alpha_{0}\rangle\otimes|e^{i\phi_{1}}\sqrt{1-\gamma_{1}}\alpha_{1}\rangle\otimes\cdots\otimes|e^{i\phi_{d}}\sqrt{1-\gamma_{d}}\alpha_{d}\rangle$ followed by measurement and estimation.
@@ III.1 Minimum QCRB of multi-mode NOON states
$\hat{\rho}_{N,\bm{\phi}}=\left(\sum_{j=0}^{d}p_{j}(1-\gamma_{j})^{N}\right)|\widetilde{\psi}_{N,\bm{\phi}}\rangle\langle\widetilde{\psi}_{N,\bm{\phi}}|+\hat{\sigma}_{N}.$ (6)
Here, $\hat{\sigma}_{N}$ is a non-negative density matrix (see Appendix A) diagonalized in the Fock state basis up to $(N-1)$, and $|\widetilde{\psi}_{N,\bm{\phi}}\rangle$ is a lossy multi-mode NOON state in the $N$-photon manifold written as
$|\widetilde{\psi}_{N,\bm{\phi}}\rangle$ $=\mathcal{N}\Bigg{(}\sqrt{p_{0}(1-\gamma_{0})^{N}}\frac{\hat{a}_{0}^{\dagger N}}{\sqrt{N!}}$
$+\sum_{j=1}^{d}\sqrt{p_{j}(1-\gamma_{j})^{N}}\frac{\hat{a}_{j}^{\dagger N}e^{iN\phi_{j}}}{\sqrt{N!}}\Bigg{)}|0\rangle^{\otimes d+1},$ (7)
with a normalization constant $\mathcal{N}$. One can analytically find the minimum QCRB over all the weighted multi-mode NOON states taking the form of Eq. (1) as
@@
For the SQL, the minimum QCRB over all multi-mode coherent states can also be found (see Appendix D for the details) and written as
@@
when $\gamma=\gamma_{j}$ $\forall j\in[1,\cdots,d]$, and
@@ III.2 Quantum advantage from two-photon NOON states
Let us here investigate a quantum advantage of the proposed scheme using the multi-mode NOON states in comparison with the SQL under the same lossy environment. The environment that can be characterized in terms of the reference mode $\gamma_{\rm ref}$ and the other modes $\gamma=\gamma_{j}$ $\forall j\in[1,\cdots,d]$, which is inspired by the distinct roles that appear through Eq. (III.1) to Eq. (10). To quantify a quantum advantage, we define the ratio of the QCRB for the considered scheme under study to that for the optimized multi-mode coherent states, written as
@@
Figure 2 illustrates the regions of photon loss rates where a quantum advantage is achieved for the scheme proposed in Ref. p.c.humphreys (see Fig. 2(a)) and for the scheme we propose with optimal multi-mode NOON states (see Fig. 2(b)) when $N=d=2$. It is evident that the colored area of Fig. 2(b) is larger than that of Fig. 2(a), which implies that a multiple-phase estimation scheme proposed in this work is more robust to the photon loss than the scheme proposed in Ref. p.c.humphreys . Also note that the proposed scheme is slightly more robust against the photon loss in the reference mode than that of the other modes.
To look into more detailed behaviors, we take two specific values of $\gamma_{\rm ref}$ from Fig. 2(b), and compare different cases including the case using balanced NOON states (i.e., $p_{j}=\frac{1}{d}$ $\forall j$). In Fig. 3(a), we show that when $\gamma=\gamma_{\rm ref}$, our scheme is equivalent to the scheme proposed in Ref. p.c.humphreys , already pointed out in Eq. (14), and the balanced NOON states are also nearly optimal due to equal losses. It can be shown that the equivalence holds for any $N$ and $d$, but the near-optimality of the balanced NOON states is degraded with increasing $\gamma$. Interestingly, on the other hand, our scheme becomes significantly advantageous when $\gamma\not=\gamma_{\rm ref}$, which is clearly shown in Fig. 3(b) for the case of $\gamma_{\rm ref}=0.5$ that precision difference between these two schemes is well-visible. It reveals that our scheme is better than the scheme in Ref. p.c.humphreys and the balanced NOON states are far from being optimal due to unequal losses. Also note that our scheme can outperform the SQL in the lossy environment, which can be observed in Fig. 3.
@@
III.3 Quantum advantage from $d$-mode $N$-photon NOON states
Here, we examine if a quantum advantage is still achievable by $d$-mode $N$-photon NOON states with $N,d>2$ and how large it is. For the purpose, we define the critical photon loss rate $\gamma_{\rm crit}$ such that a quantum advantage is observed when $\gamma_{j}\equiv\gamma\leq\gamma_{\rm crit}$ for $j\in\{1,\cdots,d\}$ where $\gamma_{\rm ref}=0.5$. That is, $\gamma_{\rm crit}$ indicates a robustness of the scheme under study. Figures 4(a) and (b) present $\gamma_{\rm crit}$ for the scheme proposed in Ref. p.c.humphreys and $\gamma_{\rm crit}$ for our scheme, respectively, when $\gamma<\gamma_{\rm ref}=0.5$ as an example that the photon loss of the reference mode is larger than that of the other modes. It is clearly shown that the scheme proposed in this work is more robust against photon loss than Humphreys et al.’s scheme proposed in Ref. p.c.humphreys . Interestingly, the proposed scheme becomes more robust as the number of phases or modes increase. This means that our scheme is useful when the number of phases to be estimated is many, and offers robust multiple-phase estimation schemes even when $N<d$.
@@ IV Quantum advantage in terms of the CRB
Until the previous section, we have demonstrated that the QCRB of an optimal multi-mode NOON state can surpass the SQL. Exploiting the argument written in Appendix A together with Ref. l.pezze , it is easy to show that a measurement achieving the QCRB has an entanglement-based projection $|\widetilde{\psi}_{N,\bm{\phi}}\rangle\langle\widetilde{\psi}_{N,\bm{\phi}}|$. However, it is known that such the measurement is experimentally challenging to implement in general s.hong2 ; j.urrehman . Thus, for experimental feasibility, we need to verify whether an estimation scheme using a measurement with a linear optical structure can provides the SQL.
@@ V Conclusion
We have proposed and studied an optimal scheme of multiple-phase estimation using the weighted multi-mode NOON states under a lossy environment. We have shown that the proposed scheme surpasses the SQL even in the presence of loss, and is more robust to loss than the scheme proposed in Ref. p.c.humphreys that employs only a particular type of multi-mode NOON state. These behaviors persist when the numbers of modes and photons increase, and especially a quantum advantage even increases with the number of modes when the reference mode is lossier than the signal modes carrying the multi-mode NOON states. When the proposed scheme is equipped with a measurement setting configurable by using a multi-mode beam-splitter and PNRDs, a quantum advantage is still achieved although sub-optimal, which provides a practical means in multiple-phase estimation under a lossy environment. We believe the results obtained and the methodology used in this work will motivate various relevant studies in robust quantum metrology under lossy and noisy environments.

### v2

@@ I Introduction
In this work, we propose an optimal multiple-phase estimation scheme using the weighted multi-mode NOON states in the presence of photon loss. Instead of actively correcting errors, which requires more resources, we rather optimize the probe states to minimize the impact of photon loss in our scheme. Analytically deriving the estimation precision for the proposed scheme, we investigate the robustness of the scheme against photon loss and compare it with the scheme using a particular type of the multi-mode NOON state proposed in Ref. p.c.humphreys , which is known to be optimal in the lossless case. A quantum advantage is also studied in comparison with the SQL set by a scheme using the weighted multi-mode pure coherent states. As a result, we find that our scheme is not only more robust against photon loss than the scheme in Ref. p.c.humphreys , but also exhibits a quantum advantage over the SQL even in the presence of loss. These findings are elaborated in more details for the cases of three- and four-mode two-photon NOON states that have been experimentally demonstrated s.hong ; s.hong2 , followed by the generalization of the scheme with increasing the numbers of modes and photons. The proposed schemes are also investigated with a measurement scheme with general structure consisting of a multi-mode beam-splitter and multiple photon-number-resolving detectors (PNRDs), offering a practical means to achieve the quantum advantage in multiple-phase estimation under a lossy environment. We expect that our results and methodology can be used for various studies in robust quantum metrology, given that loss and decoherence are inevitable.
@@ II Multiple-phase estimation scheme under lossy environment
Figure 1: (a) Illustrative representation of the relative phases $\phi_{j}=\widetilde{\phi}_{j}-\widetilde{\phi}_{0}$ and (b) multiple-phase estimation using a $d+1$-mode probe state for estimating $d$ unknown relative phases $\{\phi_{1},\cdots,\phi_{d}\}$ with a measurement. Without loss of generality, the mode 0 is considered as a reference. It is assumed that photon loss rate in the mode $j$ is $\gamma_{j}\in[0,1]$.
@@ II.1 Scheme
Let us consider a scheme for simultaneously estimating $d$ unknown relative phases $\bm{\phi}=\{\phi_{1},\cdots,\phi_{d}\}$ with $\phi_{j}=\widetilde{\phi}_{j}-\widetilde{\phi}_{0}$ in the presence of photon loss as shown in Fig. 1(a). There, we assume that the mode 0 is used as a reference upon which the phases of all the other modes are defined. We note that our estimation problem is equivalent to what was originally considered in Ref. p.c.humphreys . For the multiple-phase estimation, we employ a ($d+1$)-mode NOON state probe written by
@@
where $\hat{a}_{j}^{\dagger}$ is the photon creation operator in the mode $j$ and $p_{j}$ with $\sum_{j=0}^{d}p_{j}=1$ and $p_{j}\in[0,1]$ is the weight of the mode $j$. This probe state undergoes $d$ individual phase shifts described by a unitary operator $\hat{U}_{\bm{\widetilde{\phi}}}=\bigotimes_{j=0}^{d}\exp(i\widetilde{\phi}_{j}\hat{a}_{j}^{\dagger}\hat{a}_{j})$ with a tuple $\bm{\widetilde{\phi}}=\{\widetilde{\phi}_{0},\cdots,\widetilde{\phi}_{d}\}$ as illustrated in Fig. 1(b). For practical relevance, we take into account photon loss that could occur in the transmission channels or inefficient detectors. The photon loss can be modeled by a fictitious beam-splitter represented by the mode transformation $\hat{a}_{j}^{\dagger}\rightarrow\sqrt{1-\gamma_{j}}\hat{a}_{j}^{\dagger}+\sqrt{\gamma_{j}}\hat{e}_{j}^{\dagger}$ p.kok with photon loss rate $\gamma_{j}\in[0,1]$, where $\hat{e}_{j}^{\dagger}$ is a virtual vacuum input mode for describing environment. It can be shown, by the Kraus representation of photon loss, that photon losses occurring before and after the phase shifters are equivalent. This allows to consider only the photon loss taking place onto the phase-encoded probe state $|\psi_{N,\bm{\widetilde{\phi}}}\rangle=\hat{U}_{\bm{\widetilde{\phi}}}|\psi_{N}\rangle$ before measurement. The phase-encoded ($d+1$)-mode NOON state is thus transformed to the output state
$\hat{\rho}_{N,\bm{\widetilde{\phi}}}=\Lambda_{\gamma_{0}}\otimes\cdots\otimes\Lambda_{\gamma_{d}}(|\psi_{N,\bm{\widetilde{\phi}}}\rangle\langle\psi_{N,{\bm{\widetilde{\phi}}}}|),$ (2)
where $\Lambda_{\gamma_{j}}$ is a photon loss channel in the mode $j$ m.grassl . If $\gamma_{j}$ is zero, then the quantum channel $\Lambda_{\gamma_{j}}$ becomes an identity channel. Finally, the estimation of $d$ unknown relative phases is made based on the measurement results obtained by a measurement performed onto the output state of Eq. (2).
@@ III.1 Minimum QCRB of multi-mode NOON states
$\hat{\rho}_{N,\bm{\widetilde{\phi}}}=\left(\sum_{j=0}^{d}p_{j}(1-\gamma_{j})^{N}\right)|{\psi}_{N,\bm{\widetilde{\phi}}}^{(\bm{\gamma})}\rangle\langle{\psi}_{N,\bm{\widetilde{\phi}}}^{(\bm{\gamma})}|+\hat{\sigma}_{N}.$ (6)
Here, $\hat{\sigma}_{N}$ is a phase-independent non-negative density matrix (see Appendix A) diagonalized in the Fock state basis up to $(N-1)$, and $|{\psi}_{N,\bm{\widetilde{\phi}}}\rangle$ is a lossy multi-mode NOON state in the $N$-photon manifold written as
$|{\psi}_{N,\bm{\widetilde{\phi}}}^{(\bm{\gamma})}\rangle$ $=$ $\mathcal{N}\Bigg{(}\sqrt{p_{0}(1-\gamma_{0})^{N}}\frac{\hat{a}_{0}^{\dagger N}}{\sqrt{N!}}$ (7)
$+$ $\sum_{j=1}^{d}\sqrt{p_{j}(1-\gamma_{j})^{N}}\frac{\hat{a}_{j}^{\dagger N}e^{iN{\phi}_{j}}}{\sqrt{N!}}\Bigg{)}|0\rangle^{\otimes d+1},$
with a normalization constant $\mathcal{N}$ and relative phases $\phi_{j}=\widetilde{\phi}_{j}-\widetilde{\phi}_{0}$. By exploiting that $|{\psi}_{N,\bm{\widetilde{\phi}}}^{(\bm{\gamma})}\rangle$ in Eq. (7) is orthogonal to support of $\hat{\sigma}_{N}$ as explained in Appendix A, the QFIM $F_{\rm Q}(\bm{\phi})$ is formulated as (see Appendix B for details)
$F_{\mathrm{Q}(j,k)}(\bm{\phi})=4\left(\sum_{j=0}^{d}p_{j}(1-\gamma_{j})^{N}\right)\times$
$\left\{\langle\partial_{\phi_{j}}\psi_{N,\bm{\widetilde{\phi}}}^{(\bm{\gamma})}|\partial_{\phi_{k}}\psi_{N,\bm{\widetilde{\phi}}}^{(\bm{\gamma})}\rangle-\langle\partial_{\phi_{j}}\psi_{N,\bm{\widetilde{\phi}}}^{(\bm{\gamma})}|\psi_{N,\bm{\widetilde{\phi}}}^{(\bm{\gamma})}\rangle\langle\psi_{N,\bm{\widetilde{\phi}}}^{(\bm{\gamma})}|\partial_{\phi_{k}}\psi_{N,\bm{\widetilde{\phi}}}^{(\bm{\gamma})}\rangle\right\}.$ (8)
It can be shown that the QFIM written above is invertible, which means that every relative phases are simultaneously estimated even in the presence of photon loss. The inverse of the QFIM is analytically derived by using the Sherman-Morrison formula sherman-morrison , and the QCRB is evaluated as
$\mathrm{Tr}\left\{{F}_{\rm Q}^{-1}(\bm{\phi})\right\}$
$=\frac{1}{4N^{2}}\left\{\frac{d}{p_{0}(1-\gamma_{0})^{N}}+\sum_{j=1}^{d}\frac{1}{p_{j}(1-\gamma_{j})^{N}}\right\}.$ (9)
The QCRB written above is a convex function of $\bm{p}=(p_{0},\cdots,p_{d})$. Thus, from the fact that a point vanishing gradient of the QCRB in Eq. (III.1) is a global optimum s.boyd , one can analytically find the minimum QCRB over all the weighted multi-mode NOON states taking the form of Eq. (1) as
@@
As a classical benchmark, we consider the scheme that employs a ($d+1$)-mode coherent state described by $\otimes_{j=0}^{d}|\alpha_{j}\rangle$ with $\alpha_{j}=\sqrt{q_{j}N}$ ($q_{0}+\cdots+q_{j}=1$, $q_{j}\in[0,1]$) under the same environment r.j.glauber , leading to the output state $\otimes_{j=0}^{d}|e^{i\widetilde{\phi}_{j}}\alpha_{j}\sqrt{1-\gamma_{j}}\rangle$ followed by measurement and estimation. We define the SQL as the minimum QCRB over all multi-mode pure coherent states optimized over $q_{j}$ and (see Appendix D for the details)
@@
when $\gamma=\gamma_{j}$ $\forall j\in\{1,\cdots,d\}$, and
@@ III.2 Quantum advantage from two-photon NOON states
Let us here investigate a quantum advantage of the proposed scheme using the multi-mode NOON states in comparison with the SQL under the same lossy environment. The environment that can be characterized in terms of the reference mode $\gamma_{\rm ref}$ and the other modes $\gamma=\gamma_{j}$ $\forall j\in[1,\cdots,d]$, which is inspired by the distinct roles that appear through Eq. (III.1) to Eq. (12). To quantify a quantum advantage, we define the ratio of the QCRB for the considered scheme under study to that for the optimized multi-mode pure coherent states, written as
@@
Figure 2 illustrates the regions of photon loss rates where a quantum advantage is achieved for the scheme proposed in Ref. p.c.humphreys (see Fig. 2(a)) and for the scheme we propose with optimal multi-mode NOON states (see Fig. 2(b)) when $N=d=2$. It is evident that the colored area of Fig. 2(b) is larger than that of Fig. 2(a), which implies that a multiple-phase estimation scheme proposed in this work is more robust to the photon loss than the scheme proposed in Ref. p.c.humphreys . This is simply because our schemes employ the weighted NOON states whose weights are optimized for individual loss rates. Also note that the proposed scheme is slightly more robust against the photon loss in the reference mode than that of the other modes.
To look into more detailed behaviors, we take two specific values of $\gamma_{\rm ref}$ from Fig. 2(b), and compare different cases including the case using balanced NOON states (i.e., $p_{j}=\frac{1}{d}$ $\forall j$). In Fig. 3(a), we show that when $\gamma=\gamma_{\rm ref}$, our scheme is equivalent to the scheme proposed in Ref. p.c.humphreys , already pointed out in Eq. (16), and the balanced NOON states are also nearly optimal due to equal losses. It can be shown that the equivalence holds for any $N$ and $d$, but the near-optimality of the balanced NOON states is degraded with increasing $\gamma$. Interestingly, on the other hand, our scheme becomes significantly advantageous when $\gamma\not=\gamma_{\rm ref}$, which is clearly shown in Fig. 3(b) for the case of $\gamma_{\rm ref}=0.5$ that precision difference between these two schemes is well-visible. It reveals that our scheme is better than the scheme in Ref. p.c.humphreys when the loss rates are unequal to each other and the balanced NOON states are far from being optimal due to unequal losses. Also note that our scheme can outperform the SQL in the lossy environment, which can be observed in Fig. 3.
@@
III.3 Quantum advantage from $d$-mode $N$-photon NOON states
@@
Here, we examine if a quantum advantage is still achievable by $d$-mode $N$-photon NOON states with $N,d>2$ and how large it is. For the purpose, we define the critical photon loss rate $\gamma_{\rm crit}$ such that a quantum advantage is observed when $\gamma_{j}\equiv\gamma\leq\gamma_{\rm crit}$ for $j\in\{1,\cdots,d\}$ where $\gamma_{\rm ref}=0.5$. That is, $\gamma_{\rm crit}$ indicates a robustness of the scheme under study. Figures 4(a) and (b) present $\gamma_{\rm crit}$ for the scheme proposed in Ref. p.c.humphreys and $\gamma_{\rm crit}$ for our scheme, respectively, when $\gamma<\gamma_{\rm ref}=0.5$ as an example that the photon loss of the reference mode is larger than that of the other modes. It is clearly shown that the scheme proposed in this work is more robust against photon loss than Humphreys et al.’s scheme proposed in Ref. p.c.humphreys . Interestingly, the proposed scheme becomes more robust as the number of phases or modes increase. This means that our scheme is useful when the number of phases to be estimated is many, and offers robust multiple-phase estimation schemes even when $N<d$. Particularly for cases where $N=2$ and $d=10$, we also study general loss scenarios with arbitrary loss rates of $\gamma_{j}$’s. This involves simulating $10^{5}$ instances, where each instance’s loss rates are arbitrarily generated from a uniform distribution within a specific range $[\gamma_{\rm min},\gamma_{\rm max}]$. We find that our schemes generally perform well because our NOON states are optimized for individual loss scenarios (see Appendix F for the details).
@@ IV Quantum advantage in terms of the CRB
Until the previous section, we have demonstrated that the QCRB of an optimal multi-mode NOON state can surpass the SQL. Exploiting the argument written in Appendix A together with Ref. l.pezze , it is easy to show that a measurement achieving the QCRB has an entanglement-based projection $|{\psi}_{N,\bm{\phi}}^{(\bm{\gamma})}\rangle\langle{\psi}_{N,\bm{\phi}}^{(\bm{\gamma})}|$. However, it is known that such the measurement is experimentally challenging to implement in general s.hong2 ; j.urrehman . Thus, for experimental feasibility, we need to verify whether an estimation scheme using a measurement with a linear optical structure can provides the SQL.
@@ V Conclusion
We have proposed and studied an optimal scheme of multiple-phase estimation using the weighted multi-mode NOON states under a lossy environment. We have shown that the proposed scheme surpasses the SQL even in the presence of loss, and is more robust to loss than the scheme proposed in Ref. p.c.humphreys that employs only a particular type of multi-mode NOON state. These behaviors persist when the numbers of modes and photons increase, and especially a quantum advantage even increases with the number of modes when the reference mode is lossier than the signal modes carrying the multi-mode NOON states. When the proposed scheme is equipped with a measurement setting configurable by using a multi-mode beam-splitter and PNRDs, a quantum advantage is still achieved although sub-optimal, which provides a practical means in multiple-phase estimation under a lossy environment. The sub-optimality may be due to that eigenvectors composing an SLD operator is different from bases of the measurement with optimized multi-mode beam-splitter, which will be further discussed in the future work. We believe the results obtained and the methodology used in this work will motivate various relevant studies in robust quantum metrology under lossy and noisy environments.